# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ 5.1. Definition and Composition Theorem for Pure DP
In Section 5.2, we showed that for any positive $\delta$, the concurrent $1$-sparse parallel composition of CMs with the privacy parameter $(0,\delta)$ fails to preserve privacy. In our counterexample, the adversary leveraged the first outcome of a $(0,\delta)$-DP CM to decide whether to spend its budget on issuing non-identical queries to that mechanism. Intuitively, when $\delta>0$, there might exist certain “bad first answers” that occur with low probability; however, once we condition on their occurrences, the distribution of the second answer ceases to be differentially private.
@@
A verification function $f$ is first-pair consistent if the following holds: given an input message sequence $m_{1},\dots,m_{t}$, $f$ returns $\bot$ if $m_{1}$ is a pair of identical messages, i.e., $m_{1}=(m^{\prime},m^{\prime})$ for some $m^{\prime}$, while there exists $2\leq j\leq t$ such that $m_{j}$ is not a pair of identical messages.
@@
To formally define the restricted variant of the concurrent parallel composition of CMs, we introduce the verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$, where $FPC$ stands for first-pair consistency.
Definition 5.4 ($f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$).
For any input message sequence $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$ returns $\top$ if and only if both of the following conditions hold:
@@
We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$-concurrent composition of CMs as the FPC concurrent $k$-sparse parallel composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i}\right)_{i=1}^{k}$.
@@
For $\epsilon,\epsilon_{1},\dots,\epsilon_{k}\geq 0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the FPC concurrent $k$-sparse parallel composition of CMs with the privacy parameters $(\epsilon_{1},0),\dots,(\epsilon_{k},0)$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
@@
We denote $f$ as shorthand for $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,FPC}$. This proof is identical to the proof of Theorem 5.2, except for the following modification in the definition of $\mathcal{P}$: In the proof of Theorem 5.2, each $\mathcal{T}_{j}$ behaved identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ if $\delta_{j}^{\prime}>0$ and was defined according to Lemma 4.7 otherwise. In this proof, each $\mathcal{T}_{j}$ behaves identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ if the first message of the form $\left((q_{0},j),(q_{1},j)\right)$ satisfies $q_{0}=q_{1}$, and it is defined by Lemma 4.6 otherwise.
To implement this, $\mathcal{P}$’s state must include an additional sequence $\sigma_{3}$, storing a flag for each created mechanism, indicating whether the mechanism has previously received any messages. When $\mathcal{P}$ receives the creation query pair $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ from the verifier, it adds the pair $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime})$ to $\sigma_{2}$, appends a 00 to the end of $\sigma_{3}$, and returns $\top$.
Upon receiving a message of the form $\left((q_{0},j),(q_{1},j)\right)$, if the $j$-th element of $\sigma_{3}$ is 00, $\mathcal{P}$ sets this element to $1$ and replaces the entry $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime})$ in $\sigma_{2}$ with $(\mathcal{T}_{j},t_{j})$, defined as follows:
  • If $q_{0}=q_{1}$, then $\mathcal{T}_{j}$ is an IPM with the single initial state $s_{j}^{\prime}$ that never interacts with its right mechanism and interacts with its left mechanism exactly as $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
@@
$\mathcal{T}_{j}$ with the first definition never requires a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, while a $\mathcal{T}_{j}$ with the second definition may require at most one sample (see Remark 4.8). We will show that the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to $\mathcal{T}_{j}$s with the second definition form a sub-multiset of $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$:
By Definition 5.4 (ii), we know that for each index $j$, if $q_{0}=q_{1}$ in the first message of the form $\left((q_{0},j),(q_{1},j)\right)$, then $q_{0}=q_{1}$ in all subsequent messages of this form. Thus, by Definition 5.4 (i) and Definition 5.1 (ii), the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to $\mathcal{T}_{j}$s with the second definition (which receive at least one pair of non-identical queries) form a sub-multiset of $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$. ∎
@@ 5.3.1. Concurrent $k$-sparse Parallel Composition of IM
This scenario is an special case of FPC concurrent $k$-sparse parallel composition of CMs, where, instead of creating $\mathcal{M}_{i}$ and assigning a pair of initial states to it, the adversary sends $(\mathcal{M}_{i}^{\prime},\text{null},\epsilon_{i},\delta_{i},f_{\sim_{i}})^{2}$, where $\mathcal{M}_{i}^{\prime}$ is the CM simulating $\mathcal{M}_{i}$. Moreover, to ask a query $q$ from the $i$-th created IM, the adversary sends $\left((q,i),(q,i)\right)$. Since each $f_{\sim_{i}}$ is first-pair consistent, Theorem 5.5 applies to the concurrent $k$-sparse parallel composition of IM with privacy parameters $\left((\epsilon_{i},\delta_{i}\right)_{i=1}^{k}$ (that are differentially private w.r.t. neighbor relations).
NIMs are a special subclass of IMs that halt the communication upon receiving a second input message. Since Theorem 5.5 applies to all IMs, it also holds for the concurrent $k$-sparse parallel composition of NIMs with privacy parameters $\left((\epsilon_{1},\delta_{i}\right)_{i=1}^{k}$. In the next section, we compare this result to the classical parallel composition theorem.
@@
It is known that the $k$-sparse parallel composition of $\ell$ $(\epsilon_{0},\delta_{0})$-DP NIMs is $(\epsilon,\delta)$-DP if the composition of $\ell$ instances of $\mathrm{RR}_{\epsilon_{0},\delta_{0}}$ is $(\epsilon,\delta)$-DP. We showed that any NIM that is $(\epsilon_{0},\delta_{0})$-DP w.r.t. a neighbor relation $\sim$ can be modeled by a CM that is $(\epsilon_{0},\delta_{0})$-DP w.r.t. $f_{\sim}$, which is a first-pair consistent verification function. Thus, setting $\epsilon_{1}=\dots=\epsilon_{k}=\epsilon_{0}$, $\delta_{1}=\dots=\delta_{k}=\delta_{0}$ in Theorem 5.5 recovers and generalizes the classic parallel composition theorem. Our theorem extends differential privacy guarantees to settings where an adversary can adaptively create an unbounded number of NIMs.
@@ 6. Designing Complex Continual Mechanisms
In Section 6.1, we formalize the above properties for $\mathcal{P}$, and in Section 6.2, we prove the desired privacy guarantees for $\mathcal{M}$. Combining this result with the concurrent composition theorems provided in the previous sections, we can avoid complicated privacy analysis for many CMs.
@@
To formally state that $\mathcal{P}$ maps $f$-valid sequences to $f^{\prime}$-valid ones, we assume that $\mathcal{P}$ is deterministic. This means $\mathcal{P}$ can be expressed as a deterministic function that maps the sequence of received messages to its next output message. This assumption is not restrictive, as randomness can be modeled by a 00-DP sub-mechanism $\mathcal{M}_{i}$ that accepts a single message $\top$, and upon receiving it, generates a sufficiently long sequence of random bits or samples from a fixed distribution.
Recall that the goal of this section is to show $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $f^{\prime}$ if ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f$. That is, we need to show that the view of any adversary in the interactions with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
@@
In Theorem 6.1, we assume that the view of any adversary interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable. The goal is to show the same result for $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$. The key idea of the proof is to show that the IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{T}\!\circ^{*}\!\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are identical, where $\mathcal{T}$ is an IPM with a singleton initial state space. The proof consists of two steps.
@@
$\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}^{2}$ in Claim 6.4 is an IPM with a singleton initial state space. Thus, by Lemma 2.15, that $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}\!\circ^{*}\!{\textrm{ExtConComp}}$, is $(\epsilon,\delta)$-DP w.r.t. $\sim_{01}$ implies the view of any adversary in the interactions with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable, which finishes the proof. ∎

### v3

@@ 5.2. Composition Theorem
In Theorem 5.2, we showed that for any positive $\delta$, the concurrent $1$-sparse parallel composition of CMs with the privacy parameter $(0,\delta)$ fails to preserve privacy. In our counterexample, the adversary leveraged the first outcome of a $(0,\delta)$-DP CM to decide whether to spend its budget on issuing non-identical queries to that mechanism. Intuitively, when $\delta>0$, there might exist certain “bad first answers” that occur with low probability; however, once we condition on their occurrences, the distribution of the second answer ceases to be differentially private.
@@
A verification function $f$ is first-pair consistent if the following holds: for every input message sequence $m_{1},\dots,m_{t}$, $f(m_{1},\dots,m_{t})=\bot$ whenever $m_{1}$ is a pair of identical messages, i.e., $m_{1}=(m^{\prime},m^{\prime})$ for some $m^{\prime}$, and there exists $2\leq j\leq t$ such that $m_{j}$ is not a pair of identical messages.
@@
To formally define the restricted variant of the concurrent parallel composition of CMs, we introduce the verification function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$, where $FPC$ stands for first-pair consistency.
Definition 5.14 ($f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$).
For any $t\in\mathbb{N}$ and any input message sequence $m_{1},\dots,m_{t}$, the function $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$ returns $\top$ if and only if both of the following conditions hold:
@@
We refer to the $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$-concurrent composition of CMs as the FPC concurrent $k$-sparse parallel composition of CMs with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$.
@@
For $\epsilon_{1},\dots,\epsilon_{k}>0$, $\epsilon\geq 0$ and $0\leq\delta,\delta_{1},\dots,\delta_{k}\leq 1$, if the composition of non-interactive mechanisms $\mathrm{RR}_{\epsilon_{1},\delta_{1}},\dots,\mathrm{RR}_{\epsilon_{k},\delta_{k}}$ is $(\epsilon,\delta)$-DP, then the FPC concurrent $k$-sparse parallel composition of CMs with the privacy parameters $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$ is also $(\epsilon,\delta)$-DP against an adaptive adversary.
@@
We denote $f$ as shorthand for $f^{\epsilon_{1},\delta_{1},\dots,\epsilon_{k},\delta_{k}}_{\infty,\mathit{FPC}}$. The proof is identical to the proof of Theorem 4.11, except for the definition of IPMs $\mathcal{T}_{j}$ that are executed by the IPM $\mathcal{P}$: In the proof of Theorem 4.11, each $\mathcal{T}_{j}$ was defined according to Lemma 4.6. In this proof, each $\mathcal{T}_{j}$ is defined by Corollary 5.7 if the first message of the form $\left((q_{0},j),(q_{1},j)\right)$ satisfies $q_{0}\neq q_{1}$, and it behaves identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ otherwise.
To implement this, the sequence $\sigma_{2}$ in $\mathcal{P}$’s state must include an additional boolean variable $u_{j}$ for each created mechanism $\mathcal{M}_{j}^{\prime}$, indicating whether the mechanism has previously received any messages. Specifically, when $\mathcal{P}$ receives the creation query pair $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},f_{j}^{\prime})^{2}$ from the verifier, instead of adding $(\mathcal{T}_{j},t_{j},\epsilon_{j}^{\prime},\delta_{j}^{\prime})$, it adds the tuple $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},u_{j}=0)$ to $\sigma_{2}$.
Upon receiving a left message of the form $\left((q_{0},j),(q_{1},j)\right)$, if $u_{j}=0$, before responding, $\mathcal{P}$ replaces the $j$-th entry $(\mathcal{M}_{j}^{\prime},s_{j}^{\prime},\epsilon_{j}^{\prime},\delta_{j}^{\prime},u_{j}=0)$ in $\sigma_{2}$ with $(\mathcal{T}_{j},t_{j},\epsilon_{j}^{\prime},\delta_{j}^{\prime},1)$, where $T_{j}$ is an IPM defined as follows and $t_{j}$ is its initial state:
  • If $q_{0}=q_{1}$, then $\mathcal{T}_{j}$ is an IPM with the single initial state $\left((),0,s_{j}^{\prime}\right)$ that never interacts with its right mechanism and interacts with its left mechanism exactly as $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$. That is, upon receiving a right message, $\mathcal{T}_{j}$ halts the communication, and upon receiving a left message, it updates its state and responds identically to $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$.
@@
To generate a response to a left message of the form $\left((q_{0},j),(q_{1},j)\right)$, as before, $\mathcal{P}$ forwards $(q_{0},q_{1})$ to $\mathcal{T}_{j}$ as a left message and provide a properly chosen $r_{i}$ as a right response if needed (see the proof of Theorem 4.11).
By Definition 5.14 (ii), we know that for each index $j$, if $q_{0}=q_{1}$ in the first message of the form $\left((q_{0},j),(q_{1},j)\right)$, then $q_{0}=q_{1}$ in all subsequent messages of this form. In this case, by the above definition, for each $b\in\{0,1\}$, the mechanism $\mathcal{T}_{j}$ never interacts with $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}(b)$ and simulates $\mathcal{V}[f_{j}^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{M}_{j}^{\prime}$ identically. Moreover, when $q_{0}\neq q_{1}$ in the first message, by Definition 5.14 (i) and Definition 5.1 (ii), the privacy parameters $(\epsilon_{j}^{\prime},\delta_{j}^{\prime})$ corresponding to $\mathcal{T}_{j}$s with the second definition form a sub-multiset of $(\epsilon_{1},\delta_{1}),\dots,(\epsilon_{k},\delta_{k})$. Therefore, whenever an IPM $\mathcal{T}_{j}$ requires a sample from $\mathrm{RR}_{\epsilon_{j}^{\prime},\delta_{j}^{\prime}}$, $\mathcal{P}$ can find such sample in $\sigma_{1}$ that is unused. ∎
@@ 5.3.1. Concurrent $k$-sparse Parallel Composition of IM
This scenario is an special case of FPC concurrent $k$-sparse parallel composition of CMs, where, instead of creating $\mathcal{M}_{i}$ and assigning a pair of initial states to it, the adversary sends $(\mathcal{M}_{i}^{\prime},\text{null},\epsilon_{i},\delta_{i},f_{\sim_{i}})^{2}$, where $\mathcal{M}_{i}^{\prime}$ is the CM simulating $\mathcal{M}_{i}$. Moreover, to ask a query $q$ from the $i$-th created IM, the adversary sends $\left((q,i),(q,i)\right)$. Since each $f_{\sim_{i}}$ is first-pair consistent, Theorem 5.15 applies to the concurrent $k$-sparse parallel composition of IM with privacy parameters $\left((\epsilon_{i},\delta_{i})\right)_{i=1}^{k}$ (that are differentially private w.r.t. neighbor relations).
NIMs are a special subclass of IMs that halt the communication upon receiving a second input message. Since Theorem 5.15 applies to all IMs, it also holds for the concurrent $k$-sparse parallel composition of NIMs with privacy parameters $\left((\epsilon_{1},\delta_{i}\right)_{i=1}^{k}$. We note that the restricted version of Theorem 5.15 for the $k$-sparse parallel composition of (unbounded number of) NIMs was previously proved in [20] as an intermediate lemma. In the next section, we compare the $k$-sparse parallel composition of IMs to the classical parallel composition theorem for NIMs.
@@
It is known that the $k$-sparse parallel composition of $\ell$ $(\epsilon_{0},\delta_{0})$-DP NIMs is $(\epsilon,\delta)$-DP if the composition of $\ell$ instances of $\mathrm{RR}_{\epsilon_{0},\delta_{0}}$ is $(\epsilon,\delta)$-DP. We showed that any NIM that is $(\epsilon_{0},\delta_{0})$-DP w.r.t. a neighbor relation $\sim$ can be modeled by a CM that is $(\epsilon_{0},\delta_{0})$-DP w.r.t. $f_{\sim}$, which is a first-pair consistent verification function (see the example below Definition 5.13). Thus, setting $\epsilon_{1}=\dots=\epsilon_{k}=\epsilon_{0}$, $\delta_{1}=\dots=\delta_{k}=\delta_{0}$ in Theorem 5.15 recovers and generalizes the classic parallel composition theorem. Our theorem extends differential privacy guarantees to settings where an adversary can adaptively create an unbounded number of NIMs.
@@ 6. Designing Complex Continual Mechanisms
In Section 6.1, we formalize the above properties for $\mathcal{P}$, and in Section 6.2, we prove the desired privacy guarantees for $\mathcal{M}$. Combining this result with the concurrent composition theorems provided in the previous sections, we can avoid complicated privacy analysis for many CMs. To obtain rigorous results, our statements and proofs must be presented in a technical and somewhat complex form. In contrast, mechanisms are usually described more simply using pseudocode rather than message-passing state machines. In Section 6.3, we reformulate our results in terms of pseudocode, enabling new mechanisms to be analyzed directly without the need for reformulating to the setting of in this paper.
@@
To formally state that $\mathcal{P}$ maps $f$-valid sequences to $f^{\prime}$-valid ones, we assume that $\mathcal{P}$ is deterministic. This means $\mathcal{P}$ can be expressed as a deterministic function that maps the sequence of received messages to its next output message. This assumption is not restrictive, as randomness can be modeled by a 0-DP sub-mechanism $\mathcal{M}_{i}$ that accepts a single message $\top$, and upon receiving it, generates a sufficiently long sequence of random bits or samples from a fixed distribution.
Recall that the goal of this section is to show $\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-DP w.r.t. $f^{\prime}$ if ExtConComp is $(\epsilon,\delta)$-DP w.r.t. $f$. That is, we need to show that the view of every adversary interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable.
@@
In Theorem 6.1, we assume that the view of every adversary interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable. The goal is to show the same result for $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$. The key idea of the proof is to show that the IMs $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{T}\!\circ^{*}\!\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ are identical, where $\mathcal{T}$ is an IPM with a singleton initial state space. The proof consists of two steps.
@@
$\mathcal{V}[f]\!\circ^{*}\!\mathcal{P}^{2}$ in Claim 6.4 is an IPM with a singleton initial state space. Thus, by Lemma 2.16, that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f^{\prime}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable implies that for every adversary $\mathcal{A}$, the views of $\mathcal{A}$ interacting with $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ and $\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{P}\!\circ^{*}\!{\textrm{ExtConComp}}$ is $(\epsilon,\delta)$-indistinguishable, which finishes the proof. ∎
6.3. Simplified Formulation for Pseudocodes
Continual mechanisms are typically not described as state machines exchanging messages. Instead, they are presented in the form of pseudocode: at each time step, the pseudocode receives an input, updates its internal variables, and produces an output. For clarity and practicality, we restate the results of this section in the pseudocode framework.
Decision vs. Auxiliary Variables.
We partition the variables of a continual mechanism $\mathcal{M}$ into two categories:
  • Decision variables: These are either outputs of submechanisms or variables that are computed solely based on decision variables. Intuitively, it follows that they are post-processing of the privatized outputs.
  • Auxiliary variables: These are all remaining variables. The values of these variables may depend on the mechanism’s inputs as well as on both decision and auxiliary variables.
Note.
For simplicity, we assume pseudocodes are deterministic as any randomness could be modeled as a submechanism call. For instance, sampling from the uniform distribution could be represented as invoking a 0-DP submechanism that outputs such a sample on a fixed input.
Properties.
For $t\in\mathbb{N}$, given two sequences $\sigma^{0}=(m_{1}^{0},\dots,m_{t}^{0})$ and $\sigma^{1}=(m_{1}^{1},\dots,m_{t}^{1})$, their combination is defined as $\big((m_{1}^{0},m_{1}^{1}),\dots,(m_{t}^{0},m_{t}^{1})\big)$. Recall the three properties from Section 6.1. We reformulate them for pseudocodes as follows:
  1. (1)
    Destination property. A pseudocode satisfies this property if (i) the decision of whether to produce an output, create a new submechanism, or interact with an existing one, (ii) the choice of the new submechanism in the second case, and (iii) the selection of the submechanism to interact with in the third case is made solely based on decision variables.
  2. (2)
    Response property. A pseudocode satisfies this property if its output is determined solely by decision variables.
  3. (3)
    $f\to f^{\prime}$ property. Let $f$ and $f^{\prime}$ be verification functions. A pseudocode satisfies this property if, for every $t\in\mathbb{N}$ and every pair of input sequences $(\sigma^{0},\sigma^{1})$ of length $t$ forming an $f$-valid combination, the following holds: Assume that all decision variables of $\mathcal{M}$ are updated identically through the end of time step $t$. For each $b\in\{0,1\}$, define $\sigma^{b}$ to denote the sequence of creation queries and inputs to submechanisms when processing $\sigma^{b}$. Here, a creation query is a tuple specifying a created submechanisms, its privacy parameters, and associated verification functions. Also, an input $m$ to the $j$-th submechanism is represented by the pair $(m,j)$. Then, the combination of $\sigma^{0}$ and $\sigma^{1}$ is required to be $f^{\prime}$-valid.
Corollary 6.5.
Let $f$ and $f^{\prime}$ be two verification functions. For $\epsilon\geq 0$ and $0\leq\delta\leq 1$, suppose the $f^{\prime}$-concurrent composition of continual mechanisms is $(\epsilon,\delta)$-DP. Then, every continual mechanism whose pseudocode satisfies the destination, response, and $f\to f^{\prime}$ properties, defined above, is $(\epsilon,\delta)$-DP w.r.t. $f$.